OpenStudy (anonymous):

Assume cos(x) + 2 sin(x) =11/5 and 6 cos(x) + sin(x) =22/5 Find the exact (numeric) value of sec(x)

myininaya (myininaya):

let u=cos(x) and v=sin(x) so we have: u+2v=11/5 6u+v=22/5 solve for u and v

OpenStudy (anonymous):

ok thank you!!

OpenStudy (anonymous):

I guess it has not answer because one equation more is there. u2+v2=1 So two variable with three equations.

OpenStudy (anonymous):